# Learning Maximal Safe Sets Using Hypernetworks for MPC-based Local Trajectory Planning in Unknown Environments

### v2

@@ V-E Single-Network Architecture Analysis
To emphasize the advantages of the proposed two-network architecture, we analyze the potential use of a single-network architecture and identify the practical difficulties that arise. Inspired by the work in [17], we consider an MLP consisting of 3 hidden layers with 512 neurons. In this case, an SDF represents environment parameters that are treated as virtual states appended to the true states of the system. Since we use an SDF of size 100x100, the joint state dimension exceeds 10.000, which is three orders of magnitude higher than the examples in [17].
From the computational efficiency aspect, integrating a neural constraint with such a large number of parameters in the optimization loop significantly increases optimization time, rendering the MPC planner impractical. For example, if we use such a model as the terminal constraint for the Dubins car and horizon length of 10 steps, it takes $\sim$1.6 s to find a solution. In comparison, the mean optimization time for our approach is $\sim$5.5 ms.
@@ VI DISCUSSION ON LIMITATIONS
Local planners are typically unable to account for occlusions and long-term consequences, which prevents our method from achieving recursive feasibility in some scenarios. For example, if an obstacle is occluded due to the limited field of view of the sensor, once it suddenly appears it can potentially result in infeasibility. Also, while avoiding an obstacle may seem equally viable from either the left or right side at a given moment, one choice could result in a blocked path later in time.
The main theoretical limitation of our method is the lack of guarantees of safety, which is a common drawback of learning-based approaches [29]. Our major attempt in this direction is the proposed RWMSE loss function used during training, which improves the estimation accuracy in the vicinity of a BRT boundary. At the same time, the technical challenge is the computational complexity of the HJ reachability which scales exponentially with the model size. To improve scalability, one can use already existing techniques in the HJ reachability community, such as state decomposition [15] or self-supervised neural solvers [16] to directly support higher-dimensional systems.
@@ VII CONCLUSIONS
We present a learning-based method to approximate the HJ value function in real-time based on a high-dimensional observation of the environment, which is not possible with the existing methods in the field. Furthermore, we propos a new MPC architecture with a neural terminal constraint that uses the approximated maximal safe set as the terminal set constraint, resulting in improved recursive feasibility and safety, especially for short prediction horizons. We also introduce a custom RWMSE loss function, which achieves better results than the standard MSE loss.

### v3

@@ V-E Single-Network Architecture Analysis
To emphasize the advantages of the proposed two-network architecture, we analyze the potential use of a single-network architecture and identify the practical difficulties that arise. Inspired by the work in [14], we consider an MLP model consisting of 3 hidden layers with 512 neurons. In this case, an SDF represents environment parameters that are treated as virtual states appended to the true states of the system. Since we use an SDF of size 100$\times$100, the joint state dimension exceeds $10\times 10^{3}$, which is three orders of magnitude higher than the examples in [14]. The total number of trainable parameters is $\sim 5.6\times 10^{6}$, which is the same order of magnitude as for the hypernetwork in our approach.
From the computational efficiency aspect, integrating a neural constraint with such a large number of parameters in the optimization loop significantly increases optimization time, rendering the MPC planner impractical. For example, using such a model as the terminal constraint for the Dubins car and horizon length of 10 steps takes $\sim$1.6 s to find a solution, while our approach needs $\sim$5.5 ms on average.
@@ VI Discussion on Limitations
The main theoretical limitation of our method is the lack of safety guarantees, which is a common drawback of learning-based approaches [29]. Our major attempt in this direction is the proposed RWMSE loss function used during training, which improves the estimation accuracy in the vicinity of a BRT boundary. To achieve more desirable safety properties of the learned model, one could employ probabilistic measures of the approximation error to form chance constraints. Also, the HJ value function poses certain mathematical properties that could be exploited and embedded into the model architecture by custom design of the main network.
From the technical perspective, the major challenge arises from the curse of dimensionality associated with HJ reachability analysis during the data generation phase. The existing numerical tools are usually not applicable to systems with more than six state dimensions. To improve scalability, one can use existing techniques such as state decomposition [12] or neural solvers [13] to generate data for higher-dimensional systems. Another direction would be to replace the supervised learning procedure with other types of learning. For example, using reinforcement learning (RL) as presented in [30] or self-supervision similar to [13] are appealing options.
@@ VII Conclusions
We present a learning-based method to approximate the HJ value function in real-time based on a high-dimensional observation of the environment, which is not possible with the existing methods in the field. Furthermore, we propose a new MPC architecture with a neural terminal constraint that uses the approximated maximal safe set as the terminal set constraint, resulting in improved recursive feasibility and safety, especially for short prediction horizons. We also introduce a custom RWMSE loss function, which achieves better results than the standard MSE loss.